# LLMs’ Reshaping of People, Processes, Products, and Society in Software Development: A Comprehensive Exploration with Early Adopters

### v1

@@ 2.1. Participants and Demographics
We recruited 16 participants from tech meetup groups in the southeastern US, the researchers’ personal contacts, and LinkedIn. Snowball sampling was also employed, where participants invited individuals from their networks to join the study (Sadler et al., 2010). The target population for this study was professional software engineers and developers. Inclusion criteria required participants to have at least two months of experience using ChatGPT or a similar LLM-based chatbot for programming in their work. Initially, our call for participants attracted a number of graduate students. Although many had prior experience as developers or had worked in development roles at their universities, we determined that this group would not fully address our research questions. As a result, we revised our recruitment materials to focus specifically on full-time developers. The demographic details of all 16 participants are provided in Table 2.
2.2. Interview Questions Formulation
@@
We structured our interview questions around four key themes in software engineering: People, Processes, Products, and Society. These themes were derived from existing literature, which highlighted the critical aspects of each category. Using this foundational framework, we formulated interview questions aimed at capturing relevant insights within each theme. An iterative approach was employed in developing these questions to ensure their clarity and relevance.
@@
Following the pilot interviews, we reviewed the data and feedback, ultimately excluding the student interviews from the preliminary analysis due to concerns about their alignment with the study’s objectives. Based on the insights gained, we refined the questions, reworded them for clarity, checked for grammatical accuracy, and ensured the wording was universally understood. The final version of the questions used for the interviews can be found in Table 1. Additionally, we reduced the interview duration to ensure the questions could be answered efficiently without sacrificing the depth of the responses.
2.3. Procedure
2.3.1. Interview
Table 1. A list of the interview questions we asked participants.
@@ 2.3.1. Interview
We invited interested participants via email and asked them to complete consent forms. Participants used a Google Calendar link to schedule their own research sessions, which lasted approximately 70 minutes. While all participants consented to audio recording, not all agreed to screen or video capture. All interviews took place between the period of March 1, 2023, to July 7, 2023.
The study consisted of semi-structured interviews, featuring a core set of predetermined questions, while allowing flexibility for follow-up questions based on the conversation (Miles and Gilbert, 2005). At the start of each session, participants were briefed on the research purpose, followed by an interview of about 60 minutes. Two researchers were present: one conducted the interview, while the other took notes. Both researchers asked clarifying or follow-up questions based on participant responses.
During the interviews, participants were asked to discuss their use of LLMs in their daily work, how they evaluated its generated code, and how they integrated it into existing codebases. We also explored their opinions on how LLMs might impact the skills and job landscape in the software industry, with specific follow-up questions regarding the skills needed to effectively use LLMs. Additionally, participants were asked to describe real projects they had worked on. A list of the pre-prepared questions is provided in Table 1. While the questions we asked explicitly named ChatGPT, we stated during the interviews that interviewees should also answer the questions based on their usage of similar LLM-based chatbot other LLM-based tools.
2.3.2. Demographic Survey
Table 2. Demographics of participants P1-P16 (P#) showing the company size (number of employees), gender, race/ethnicity, years of development experience, and LLM tool(s) used.
@@ 2.3.2. Demographic Survey
After completing the interview, participants were asked to complete a brief survey on Qualtrics, which collected demographic information. Specifically, participants were asked to provide details about the size of their company, as well as their gender, ethnicity, and race. This demographic data was gathered to explore potential differences in LLMs usage across various groups. Based on the information provided during the interviews, we also estimated the number of years of experience each participant had in the software development field. These results are presented in Table 2. Our participants were 43.75% White, 43.75% Asian, and 12.5% Black. 75% were men, 12.5% were women, and 12.5% were non-binary. 75% of our participants had at least 3 years of development experience, with only four participants having more limited experience and two participants expressing they had begun their careers within the past year.
2.4. Analysis
@@
All interviews were recorded, automatically transcribed, and then reviewed and corrected by the researchers, who rewatched the video recordings to ensure accuracy.
@@
Our analysis followed an inductive approach, similar to that of Silva et al. in their study of AR activists (Silva et al., 2022). Initially, we created open codes based on three transcripts that we identified as thematically rich and highly relevant to our research questions. After discussing and refining these codes, we compiled a preliminary codebook with descriptions for each tag. The two researchers who conducted the interviews, along with one researcher who did not participate in the interviews, independently coded the remaining 13 transcripts. They cross-referenced these codes with the interview video recordings when needed. Following independent coding, the researchers held discussions to reconcile their codes, developing a unified set of agreed-upon tags that contributed to an evolving codebook. The codebook grew iteratively as new themes emerged from the diversity of participant responses.
@@
Subsequently, three researchers (two involved in coding, one external) grouped the tags according to each research question, then clustered them into mid-level themes. These were further divided into lower-level themes, which served as the foundation for our written analysis. This process took approximately two weeks, consisting of multiple intensive coding sessions. Throughout the analysis, the researchers continued to refine the codes and themes through ongoing discussions and by sharing draft results. The final coding resulted in 361 total codes, categorized into 138 low-level themes. For each theme, we also quantified the number of quotes tagged, which we present in our results. Participant quotes are referenced as PX, where X denotes the participant’s interview order. Note that P1 was technically a pilot interview, but answered all questions that were asked of the other participants.
3. Results
In this section, we present key themes found in our analysis regarding how LLMs impact people (RQ1), processes (RQ2), products (RQ3), and the environment (RQ4). We organized this section to outline the positive and negative aspects of LLMs.
3.1. RQ1: How do LLMs affect developers?
Figure 1. Findings related to RQ1, categorized by their uses (indicated by a checkmark), existing challenges (indicated by a stop sign), and challenges mentioned by our participants that have been largely addressed in newer versions of LLMs (indicated by a blue circle).
@@ 3.1. RQ1: How do LLMs affect developers?
3.1.1. Boosting Developers’ Productivity
  • Reducing Mundane Tasks: Codes related to LLMs’ effectiveness in simplifying mundane tasks appeared in fourteen interviews, aligned with previous research describing how developers used ChatGPT to automate tedious tasks (Hörnemalm, 2023) [^4] Master’s thesis. A notable example was provided by P14, who shared, “I really like that the repetitive, boring tasks — like looking for a comma — that I don’t need to do those, and I can focus more on building things.
    Thirteen participants highlighted LLMs’ role in expediting the software engineering process and saving time. For instance, P1 noted the potential benefits of integrating ChatGPT into developers’ workflows: “It might make you aware of things that you might be missing […] maybe speed up some small pieces of a software engineer’s process.” These findings align with those of other researchers (Nikolaidis et al., 2023; Rahmaniar, 2023; Andersson and Olsson, 2023).
@@ 3.1.1. Boosting Developers’ Productivity
    Similarly, nine participants highlighted how LLMs enhance efficiency and make developers more productive. For example, P15 explained how ChatGPT writes functions, supporting focus on higher-level tasks: “[Suppose that] I want to do something simple […] I don’t write those functions anymore. I always have ChatGPT do it because I know that it’s going to come up with something close to what I was going to do, but I didn’t actually have to do it. So it’s kind of like a pair-programmer for me, so I can stick to the higher-level stuff that would take more understanding of the infrastructure.”
@@
  • Streamlining Search Experiences: Ten participants highlighted the productivity advantage of LLMs in reducing search time for solutions and drew parallels between LLMs and search engines. For example, P6 shared, “[…] instead of going out directly to Google or Stack Overflow, my first line of question is ChatGPT […] I get answers a little bit faster and more contextualized [with ChatGPT].” In another case, P14 mentioned how LLMs transformed his search style: “Now, I haven’t been on Google probably for 3 weeks […] I don’t need to click on each of the links from Google to find the solution[…] [ChatGPT] gives me a summarized version from possibly three, four different sources.” In a related finding, a study by Xu et al. revealed that individuals who used ChatGPT to search for answers spent less time than their counterparts who used Google (Xu et al., 2023)  [^5] Pre-published via arXiv..
@@
  • Providing Boilerplate Code and Templates: Ten participants shared that they had used LLMs to create generic or standardized code and templates. P4 described this capability as “one of the best things I get from [ChatGPT].” P12, a consistent user of Bard [^6] Now known as Gemini, as noted in Section 2., expressed, “You can take [Bard’s response] as an idea, like a first draft, and then iterate on it -— that’s the experience I have working with it from the coding side.” Similarly, P5 mentioned, “I would ask it particular questions regarding a specific front-end component I’m developing, and I use it to point me in the right direction with code that I can start off with. And then, based on my business need, I would customize it to fit the specific requirement at the time.
  • Translating Code: Code translation was referenced in three interviews. P5 described his experience, stating, “[If] I want [the code] to be styled in a particular way, or in a certain language, or if I want to transpile it to another language, [ChatGPT has] been able to do that pretty well.” This finding is in line with prior research on code translation as one of the capabilities of AI coding tools (Lau and Guo, 2023).
  • Accelerating Learning: Two participants noted that ChatGPT reduces the learning curve and speeds up learning. Given the integral role of continuous learning in a developer’s responsibilities, LLMs’ ability to facilitate this is highly beneficial. For instance, P14 highlighted, “I feel it’s really good at explaining. Like, I don’t understand quantum computing at all, but I ask [ChatGPT] to explain the principles to me as if I’m five [years old] or as if I’m a JavaScript developer […] It helps me understand the context without needing to delve into all the underlying physics. It definitely helps me learn new things faster.” Prior research has also highlighted the positive impact of ChatGPT on accelerating the learning process (Malinka et al., 2023).
  • Simplifying Set-ups: Two participants mentioned using LLMs for set-ups and installations, which are an inevitable part of a developer’s work and can consume significant time. In one instance, P4 stated, “It saves me the four hours of headache of setting up. It would also help with setting up a new [development] environment — I’m gonna go to React, gotta get this web app building and running. I’ll just copy and paste the terminal errors, and it does a surprisingly good job of telling me how to get through the dev environment issues.
  • Supplementing Tutorials and Documentation: One participant, P16, highlighted the common issue of online tutorials and documentation being incomplete or outdated, potentially creating roadblocks in developers’ projects. She suggested that ChatGPT could alleviate this challenge: “I’m working with a platform that I’m not familiar with, and the documentation of that platform is not clear or may skip steps. I notice now that at work, documentations do skip steps, and ChatGPT does a really good job at filling in the gaps when I ask my question: ’Hey, this is what I’m supposed to be getting, but I’m getting something else. Do you know why?’ And then again, ChatGPT sometimes does give wrong answers, but it shoots out answers that I may not have considered.
  • Improving Recall of Syntax/Implementation: One participant highlighted the challenge of forgetting syntax or implementation details in a developer’s knowledge base. P11 shared, “I used to use MySQL a long time ago, and then I haven’t used MySQL in a while, so I didn’t remember certain ways of doing things. In one case, I needed to do a left join to find things that were in one table but not another. [ChatGPT] was helpful.
@@
  • Programming Language Mix-Ups: Three participants noted instances where LLMs’ responses were not in the expected language. For instance, P8 recounted an experience of asking ChatGPT to refactor code in Java: “ When you paste your code [on ChatGPT], it even does not understand that code is [in] Java, C#, JavaScript, or whatever…Maybe the premium version doesn’t have that issue. But at least the [free] version that they have is silly in those cases.” Consequently, developers need to spend time converting the code to the target language.
  • Contradictory Answers: One participant, P1, observed that ChatGPT’s answers sometimes contradicted its earlier responses and remained stubborn. He stated, “[The impact of follow-up questions on the accuracy of the answers] depends. These language models have a tendency to be stubborn.” Additionally, he cited an example supposedly from the literature where ChatGPT’s responses to an imaginary scenario were contradictory, and it refused to correct its mistakes [^7] The scenario in question concerns asking ChatGPT what it would do if it were invisible, with contradictions arising when it responds that it would access inaccessible areas; note that the researchers were unable to find the source of this reference. Despite such occasional unreliability, prior research indicates that individuals are not deterred from using ChatGPT for this reason (Amaro et al., 2023).
  • Slow Response Generation: Different versions of LLMs may exhibit varying response times that impact the user’s experience. P16 compared GPT-3.5 and GPT-4, noting significant differences in speed that resulted in frustration: “GPT-4 is really slow and very limited, so right now, I use GPT-3.5. I am excited for GPT-4 [to become faster] because it obviously holds more information. But I tried it out, and I caught myself trying to rip off my hair because it wasn’t understanding me the way that GPT-3.5 understood me. And then, when you’re limited — I think it’s 50 prompts in four hours or something like that. And I’m having conversations with ChatGPT [that are] one thread with 20 messages. […] 50 [prompts] every four hours is not efficient.
  • Struggles with Unstructured Data Analysis: One participant, P3, discussed a challenge with unstructured input for ChatGPT. As he described, “I tried asking it to perform some analysis on unstructured data I received that I was going to run through Splunk, which is a machine data reading tool. I asked it to test efficacy. I provided it with some dummy data and asked for it to create a few reports. It seems like it only read the variable names and decided what those reports would look like, but the actual visualizations it created and the insights it provided were not accurate. This might have been due to the nature of the data, but I also noticed similar issues discussed on Reddit regarding Kaggle data analysis.” Such inaccuracies could consume developers’ time to correct or result in wasted efforts when using LLMs.
3.1.2. Facilitating Developers’ Learning
@@
  • Personalizing Education and Skill Enhancement: Learning new technologies, software, and information from LLMs emerged as a prominent theme, appearing in fourteen interviews. A key benefit of LLMs is their ability to offer personalized and interactive learning. Whether LLMs introduced unfamiliar libraries to a programmer or assisted developers in learning specific languages for job interviews, they proved to be an invaluable resource. For example, P7 highlighted the constant need to learn new things in the industry: “We still need to read a lot of things like documentations or …some new technologies. So ChatGPT is a good resource…to learn something I never heard before. So, for example, if there is a question [and] I’m not sure which tool I should use, I could probably just ask the open question from ChatGPT.
@@
  • Explaining Code: We found LLMs’ abilities to provide explanations and examples to be noted in eleven interviews. For instance, P7 highlighted how a code example from ChatGPT not only offered guidance but also inspired him to clean his code: “Instead of looking at the [Golang [^8] Go, also known as Golang, is a programming language designed by Google (n.a., 2024a).] code myself — because the source code has so many packages, I don’t know where to look into — I just asked ChatGPT to find me an example. In this case, I can quickly notice this is [what] the source code is doing. I can probably do something similar to make the code readable. I mean, it has good readability and is also as beautiful as possible.” Related to LLMs’ streamlining of the searching process, P13 shared how ChatGPT was more useful than other tools due to its ability to explain code more efficiently: “[My] first step tends to be to go to ChatGPT, give it the snippet of my code, […] and then ask it those specific questions because I think it puts you quite ahead in understanding and making progress with things.
  • Providing a Broad Scope of Knowledge and Diverse Datasets: The extensive knowledge and diverse datasets of LLMs appeared in four interviews as an element that advanced developers’ knowledge and understanding. For example, P8 shared an example in which ChatGPT was able to create data transfer objects (DTOs) based on requirements related to a task in the financial domain; the LLM was able to intuit that the DTOs would require a first name, last name, SSN, PIN number, and a card number: “That was really surprising to me that [ChatGPT] knows that.
  • Demonstrating Best Coding Practices: Two participants highlighted how LLMs helped them learn how to write code more efficiently and elegantly. By observing the generated code and incorporating LLMs’ suggestions, they were able to refine their coding style and produce more readable code. P16 shared that ChatGPT helped her learn how to write code more efficiently: “I think it showed me better […] shorthand code. […] I just kind of learned how to use less words, [and] code more efficiently. I would say sometimes it may not be pretty, but from what I know, plus what I’ve learned from it, I’m able to kind of combine it and mix it and make my code more elegant and readable and just better-looking code.
@@
  • Hallucinations and Incomplete Responses: LLMs occasionally produce erroneous or fictitious content, a major issue that has been outlined in several academic studies (Tang et al., 2023b; Alkaissi and McFarlane, 2023). This challenge surfaced in six interviews. For instance, P1 recounted instances where ChatGPT fabricated responses: “There were some things that were surprising in a bad way, like it would make up papers. It would hallucinate paper names and authors.” Another participant, P11, highlighted ChatGPT’s failure to provide sources for its responses, noting the importance of human expertise and the using reliable, up-to-date information: “[People’s] knowledge is probably more up to date. No, nobody is always 100% up-to-date, but you can trust it a little bit more, or if they have experience with the library, then I would trust a person’s experience more than whatever ChatGPT is drawing from. […] ChatGPT really does not cite its information. I think maybe Bard started citing stuff. Citations are very helpful.” Additionally, P5 mentioned encountering disrupted answers midway, possibly due to server connectivity issues: “I have run into issues where I tried to generate like a lot of code, and then it would just stop halfway through because of a slow connection, or just the plan that I have for it, and then I would ask it to finish, and then it just sort of forgets what it was doing.
  • Limited Knowledge and Datasets: Contrary to the broad knowledge discussed earlier, there were five participants who shared that LLMs have limited knowledge and datasets. For instance, P9 shared that despite extensive efforts in feeding information to ChatGPT and careful prompt engineering, ChatGPT failed to provide an answer: “There are definitely times when I spent maybe more than an hour trying to ask ChatGPT to help me debug my JavaScript code. But it turns out that no matter how well I try to re-prompt, using the prompt engineering techniques that I picked up from Dr. Andrew Ng’s course on DeepLearning.AI [^9] https://www.deeplearning.ai/, [or] even if I tried all the techniques available, it’s still going in loops. ChatGPT is still unable to pick up any specific bugs, which [are] helpful for me to overcome the issue. In that case, I still have a JavaScript subcontractor that I hire on an hourly basis to help me when I’m really, really, really stuck. So there are times that I still need to get him to fill me [in on] that.
@@
  • Struggles with Novel Ideas and Logical Prompts: In four interviews, participants highlighted LLMs’ limitations in generating code for novel ideas and handling prompts requiring complex logical reasoning. For example, as P11 observed, “ChatGPT isn’t good at logical things or counting and math, right? But I think it’s able to usually generate some reasonable code for that.
  • Impediment of Developers’ Learning: Another concern raised in four interviews was LLMs’ probable adverse effect on learning, particularly for junior developers. This was primarily due to users’ potential inability to parse the correctness of ChatGPT’s answers. For example, P10 mentioned, “I think ChatGPT is not mature enough yet for everyday uses, especially for junior developers. If they want to start to do something, I don’t think they should use ChatGPT because it might lead to misinterpretation of a lot of documentation and lead to something else. Because a lot of time, ChatGPT’s tone, the way it replies, has certainties, say, here are all the solutions […] and that is incorrect. I have encountered so many times that it’s answering itself in circles.
  • Safeguard Failures and Harmful Effects: In two interviews, participants reported incidents where these LLMs’ safeguards against disseminating harmful information could be bypassed. One participant, P3, detailed an instance of impersonation that led to offensive and discriminatory content. In another case, P4 described how he was able to write a program that caused his computer to restart:“[I]t will gladly do harmful side effects. But I had to go out of my way to try to get it to [do] it. This was just during a live demo of using GPT-3, and I had it restart my computer.” These examples underscore the need for ongoing enhancements in LLMs’ capabilities to ensure their effectiveness as a knowledge-enhancing tool.
3.1.3. Supporting Developers’ Personal Growth

### v2

@@ 3.1 Participants and Demographics
We recruited 16 participants from technology meetup groups in the Southeastern United States, as well as through the researchers’ personal contacts and LinkedIn. Snowball sampling was also employed, where participants invited individuals from their networks to join the study [57]. The target population for this study was professional software engineers and developers. Inclusion criteria required participants to: (1) be employed as full-time professional software engineers or developers; (2) have begun using ChatGPT or similar LLM-based tools between November 2022 and April 2023 (qualifying them as early adopters); and (3) have at least two months of active experience using these tools for programming tasks in their professional work by the time of the interview. Initially, our call for participants attracted a number of graduate students. Although many had prior experience as developers or had worked in development roles at their universities, we determined that this group would not fully address our research questions. As a result, we revised our recruitment materials to focus specifically on full-time developers. The demographic details of all 16 participants are provided in Table 2. Our sample size of 16 participants is consistent with recommendations for qualitative interview studies seeking thematic saturation, where Guest et al. found that 12-15 interviews typically suffice to identify the majority of themes in relatively homogeneous populations [58]. Our sample size enabled us to reach thematic saturation across our four research dimensions while maintaining the depth of analysis expected in qualitative inquiry.
3.2 Interview Questions Formulation
@@
We structured the interview protocol around four key themes in software engineering: People, Processes, Products, and Society. These themes were derived from prior work on LLMs in software engineering and sociotechnical views of software development, which emphasize people, processes, products, and societal context as critical dimensions. Using this foundational framework, we formulated interview questions designed to capture relevant insights within each theme. We developed the questions iteratively to ensure their clarity, relevance, and alignment with our research questions.
@@
Following the pilot interviews, we reviewed the data and feedback, ultimately excluding the student interviews from the preliminary analysis due to concerns about their alignment with the study’s objectives. Based on the insights gained, we refined the questions, reworded them for clarity, checked for grammatical accuracy, and ensured the wording was broadly accessible to participants with different professional backgrounds. The final version of the questions used for the interviews can be found in Table 1. Additionally, we reduced the interview duration to ensure the questions could be answered efficiently without sacrificing the depth of the responses.
3.3 Procedure
3.3.1 Interview
Table 1: A list of the interview questions we asked participants.
@@ 3.3.1 Interview
We invited interested participants via email and asked them to complete consent forms. Participants used a Google Calendar link to schedule their own research sessions, which lasted approximately 70 minutes. All interviews took place between March 1, 2023, and July 7, 2023. While all participants consented to audio recording, not all agreed to the screen or video capture. Audio recordings provided sufficient data for our analysis, as our focus was on participants’ experiences and perspectives rather than detailed observation of their interactions with tools.
The study consisted of semi-structured interviews, featuring a core set of predetermined questions, while allowing flexibility for follow-up questions based on the conversation [59]. At the start of each session, participants were briefed on the research purpose (approximately 10 minutes), followed by the semi-structured interview (approximately 60 minutes). Two researchers were present: one conducted the interview, while the other took notes. Both researchers asked clarifying or follow-up questions based on participant responses.
During the interviews, participants were asked to discuss their use of LLMs in their daily work, how they evaluated LLM-generated code, and how they integrated it into existing codebases. We also explored their opinions on how LLMs might impact the skills and job landscape in the software industry, with specific follow-up questions regarding the skills needed to effectively use LLMs. Additionally, participants were asked to describe real projects they had worked on. While the questions we asked explicitly named ChatGPT, we stated during the interviews that interviewees should also answer the questions based on their usage of similar LLM-based chatbots and other LLM-based tools.
3.3.2 Demographic Survey
Table 2: Demographics of participants P1-P16 (P#) showing the company size (number of employees), gender, race/ethnicity, years of development experience, and LLM tool(s) used.
@@ 3.3.2 Demographic Survey
After completing the interview, participants were asked to complete a brief survey on Qualtrics, which collected demographic information. Specifically, participants were asked to provide details about the size of their company, as well as their gender, ethnicity, and race. This demographic data was gathered to explore potential differences in LLM usage across various groups. Based on the information provided during the interviews, we also estimated each participant’s years of experience in software development. These results are presented in Table 2. Our participants were 43.75% White, 43.75% Asian, and 12.5% Black. 75% were men, 12.5% were women, and 12.5% were non-binary. 75% of our participants had at least 3 years of development experience, with only four participants having more limited experience and two participants expressing they had begun their careers within the past year.
3.4 Analysis
@@
All interviews were recorded, automatically transcribed, and then reviewed and corrected by the researchers, who re-watched the video recordings to ensure accuracy.
@@
Our analysis followed an inductive approach, similar to that of Silva et al. in their study of AR activists [60]. Initially, we created open codes based on three transcripts that we identified as thematically rich and highly relevant to our research questions. After discussing and refining these codes, we compiled a preliminary codebook with descriptions for each tag. The two researchers who conducted the interviews, along with one researcher who did not participate in the interviews (providing an independent perspective without potential interview bias), independently coded the remaining 13 transcripts. Following independent coding, the researchers held discussions to reconcile their codes. All tags included in the final codebook required agreement from at least two of the three coders, with disagreements resolved through discussion until consensus was reached. This collaborative approach ensured coding consistency while preserving the emergent, inductive nature of our analysis and allowed us to leverage multiple perspectives in interpreting participant experiences.
@@
Subsequently, three researchers (two involved in coding and one external) grouped the tags according to each research question and then clustered them into mid-level themes. These were further divided into lower-level themes, which served as the foundation for our written analysis. This process took approximately two weeks, consisting of multiple intensive coding sessions. Throughout the analysis, the researchers continued to refine the codes and themes through ongoing discussions and by sharing draft results. The final coding resulted in 361 total codes, categorized into 138 low-level themes. For each theme, we also quantified the number of quotes tagged, which we present in our results. Participant quotes are referenced as PX, where X denotes the participant’s interview order.
To ensure analytic rigor, we employed several strategies to mitigate researcher bias. First, we included one researcher in the coding process who had not conducted interviews, providing an external perspective. Second, we maintained detailed audit trails of our coding decisions and theme development. Third, we grounded our interpretations in direct participant quotes and quantified the prevalence of themes. Finally, we remained reflexive about our own positions as researchers in academia, studying professional practice, and carefully distinguished between participants’ perspectives and our interpretations.
4 Results
In this section, we present findings from our thematic analysis of 16 interviews with early-adopter software professionals. We organize our results around four research questions that comprehensively address the sociotechnical dimensions of LLM integration: how LLMs affect developers (RQ1), influence development processes (RQ2), impact software products (RQ3), and shape broader industry and educational contexts (RQ4). For each research question, we present both benefits and challenges identified by participants, with theme prevalence indicated by the number of interviews in which each theme emerged. All findings are grounded in participant experiences and illustrated with representative quotes.
4.1 RQ1: How do LLMs affect developers?
@@
4.1.1 Boosting Developers’ Productivity
@@
  • Reducing Mundane Tasks: Codes related to LLMs’ effectiveness in simplifying mundane tasks appeared in fourteen interviews, aligned with previous research describing how developers used ChatGPT to automate tedious tasks [61]. A notable example was provided by P14, who shared, “I really like that the repetitive, boring tasks, like looking for a comma; that I don’t need to do those, and I can focus more on building things.” Thirteen participants highlighted LLMs’ role in expediting the software engineering process and saving time, while nine participants emphasized how LLMs enhance efficiency and make developers more productive. For example, P15 explained how ChatGPT writes functions, supporting focus on higher-level tasks: “[Suppose that] I want to do something simple […] I don’t write those functions anymore. I always have ChatGPT do it because I know that it’s going to come up with something close to what I was going to do, but I didn’t actually have to do it. So it’s kind of like a pair-programmer for me, so I can stick to the higher-level stuff that would take more understanding of the infrastructure.
@@
  • Streamlining Search Experiences: Ten participants highlighted the productivity advantage of LLMs in reducing search time for solutions. For example, P14 mentioned how LLMs transformed his search style: “Now, I haven’t been on Google probably for 3 weeks […] I don’t need to click on each of the links from Google to find the solution[…] [ChatGPT] gives me a summarized version from possibly three, four different sources.
@@
  • Providing Boilerplate Code and Templates: Ten participants shared that they had used LLMs to create generic or standardized code and templates. P4 described this capability as “one of the best things I get from [ChatGPT].” P12 expressed, “You can take [Bard’s response] as an idea, like a first draft, and then iterate on it; that’s the experience I have working with it from the coding side.
  • Additional Productivity Benefits: Code translation was referenced in three interviews. Two participants noted that ChatGPT reduces the learning curve and speeds up learning. Two participants mentioned using LLMs for set-ups and installations. One participant highlighted how ChatGPT could supplement incomplete or outdated tutorials and documentation, while another noted its usefulness in improving recall of syntax and implementation details.
@@
  • Programming Language Mix-Ups: Three participants noted instances where LLMs’ responses were not in the expected language, requiring developers to spend time converting the code to the target language.
  • Other Challenges: One participant observed that ChatGPT’s answers sometimes contradicted its earlier responses. One participant noted slow response generation with GPT-4 impacted user experience. One participant discussed challenges with unstructured data analysis.
4.1.2 Facilitating Developers’ Learning
@@
  • Personalizing Education and Skill Enhancement: Learning new technologies, software, and information from LLMs emerged as a prominent theme, appearing in fourteen interviews. A key benefit of LLMs is their ability to offer personalized and interactive learning. For example, P7 highlighted the constant need to learn new things in the industry: “We still need to read a lot of things like documentations or …some new technologies. So ChatGPT is a good resource…to learn something I never heard before. So, for example, if there is a question [and] I’m not sure which tool I should use, I could probably just ask the open question from ChatGPT.
@@
  • Explaining Code: LLMs’ abilities to provide explanations and examples were noted in 11 interviews. P13 shared how ChatGPT was more useful than other tools due to its ability to explain code more efficiently: “[My] first step tends to be to go to ChatGPT, give it the snippet of my code, […] and then ask it those specific questions because I think it puts you quite ahead in understanding and making progress with things.
  • Providing a Broad Scope of Knowledge and Diverse Datasets: The extensive knowledge and diverse datasets of LLMs appeared in four interviews as an element that advanced developers’ knowledge and understanding.
  • Demonstrating Best Coding Practices: Two participants highlighted how LLMs helped them learn how to write code more efficiently and elegantly by observing the generated code and incorporating LLMs’ suggestions.
@@
  • Hallucinations and Incomplete Responses: This challenge surfaced in six interviews. For instance, P1 recounted instances where ChatGPT fabricated responses: There were some things that were surprising in a bad way, like it would make up papers. It would hallucinate paper names and authors.” Another participant, P11, highlighted ChatGPT’s failure to provide sources for its responses, noting the importance of human expertise: [People’s] knowledge is probably more up to date […] ChatGPT really does not cite its information. I think maybe Bard started citing stuff. Citations are very helpful.
  • Limited Knowledge and Datasets: Five participants shared that LLMs have limited knowledge and datasets. For instance, P9 shared that despite extensive efforts and careful prompt engineering, ChatGPT failed to provide an answer, requiring him to hire a JavaScript subcontractor.
@@
  • Struggles with Novel Ideas and Logical Prompts: In four interviews, participants highlighted LLMs’ limitations in generating code for novel ideas and handling prompts requiring complex logical reasoning.
  • Impediment of Developers’ Learning: Four interviews raised concerns about LLMs’ probable adverse effect on learning, particularly for junior developers, primarily due to users’ potential inability to parse the correctness of ChatGPT’s answers.
  • Safeguard Failures and Harmful Effects: In two interviews, participants reported incidents where LLMs’ safeguards against disseminating harmful information could be bypassed.
4.1.3 Supporting Developers’ Personal Growth